the perimeter of a college athletic field is 92 meters and the length is 12 m more than the width. find the length and width.
Answer by sophxmai(62)

the perimeter of a college athletic field is 92 meters and the length is 12 m more than the width. find the length and width.
Let x=length and y=width
2x+2y=92
x-12=y

Substitute the 2nd equation into the first one
2x+2y=92
2x+2(x-12)=92
2x+2x-24=92
4x=116
x=116/4
x=29

Substitute x into one of the original equations
x-12=y
29-12=y
17=y

the length is 29m, the width is 17m